The perimeter of a rectangle is 44cm. The area of the rectangle is 117cm2. Find the length of the shorter side of the rectangle.
Answer by orca(409)

Let the length and width of the rectangle be x and y respectively.
As the perimeter of a rectangle is 44, we have
2x + 2y = 44 ......(1)
As the area of the rectangle is 117, we have
xy = 117 ........(2)
Simplifying equation (1), we have
x + y = 22 .......(3)
Solving the following simultaneous equations for x and y.
x + y = 22
xy = 117
Substituting y = 22 - x into xy = 117,we have
x(22 - x) = 117
22x - x^2 = 117
x^2 - 22x + 117 = 0
(x - 9)(x - 13) = 0
So x = 9 or x = 13
Substituting x = 9 into y = 22 - x,we have y = 13.
Substituting x = 13 into y = 22 - x,we have y = 9.
So the dimensions of the rectangle are 9 and 13.